# MoNE: Replacing Redundant Experts with Lightweight Novices for Structured Pruning of MoE

### v1

@@ 5.1 Experiment setup
5.2 Effectiveness evaluation
This section validates the effectiveness of MoNE by comparing the zero shot performance with 100 calibration samples from the Zyda2 dataset. The results are presented in Table 1. This table indicates that MoNE consistently outperforms baseline methods in terms of the average accuracy on the nine tasks. In particular, compared to baseline methods, it improves average accuracy by 2.72 for the 25% pruned OLMoE and by 3.61 for the 50% pruned Deepseek-V2-Lite. Notably, it incurs accuracy drop as small as only 1.04 for the 25% pruned Deepseek-V2-Lite. Furthermore, MoNE-pruned models achieves either the best or the second best result for individual tasks only except for the 50% pruned Moonlight on the MMLU task.
An interesting observation is that all the three expert pruning methods, MC-SMoE, RS and MoNE can achieve results on par with or even better than the original models on certain tasks. The specific examples are shown in green background in Table 1. All these results indicate that there is indeed redundancy existing in the expert level for the three pretrained MoE models, and expert pruning can rule out such redundancy to achieve even better results on these tasks. Besides, compared to MC-SMoE and RS which keep the same activated parameters, MoNE introduces less activated parameters for tokens routed to novices. Nevertheless, MoNE consistently surpasses the two baseline methods, demonstrating its strong capability.
5.3 Robustness evaluation
This section evaluates the robustness of MoNE across three key dimensions: model architecture (OLMoE, Moonlight and Deepseek-V2-Lite), calibration data sources (Zyda2 and C4), and calibration sample sizes (100, 500 and 1000). For each dimension, we vary one factor while averaging results over the other two, measuring both average accuracy and the standard deviation of accuracy drop. The results are visualized in Figure 3, with detailed scores provided in Appendix A. As shown in Figure 3(a), MoNE advances the Pareto frontier across all three dimensions at the 25% pruning ratio, demonstrating superior robustness and effectiveness compared to existing structured pruning methods. At the 50% pruning ratio (Figure 3(b)), MoNE exhibits slightly higher variance under varying model architectures and calibration sample sizes. Nevertheless, it remains the most effective method, outperforming baseline methods by a significant margin of 2.85.
@@ 5.4 Ablation study
Figure 4: Ablation study on expert access frequency, output variance and novice replacement. Numbers are the difference to the proposed MoNE. The detailed result is provided in Appendix C.
This section presents the ablation study to evaluate the effects of the two redundancy metrics and the impact of novice replacement across the downstream tasks. Figure 4 displays the average accuracy drop relative to our proposed methods, with lower values indicating greater degradation. Results are averaged over the three evaluation dimensions to provide a robust assessment.
We observe that integrating the fused expert redundancy score with novice replacement yields better performance, particularly under higher pruning ratios. This indicates that our approach is especially effective in preserving model quality when pruning is more aggressive. Notably, for tasks such as BoolQ, COPA, and PIQA, our proposed method outperforms the ablation baselines by a large margin—achieving accuracy gains of up to 8.85. However, for MMLU, pruning based solely on frequency appears to offer a slight advantage, suggesting that frequently activated experts may play a more critical role in domain-specific reasoning tasks.
@@ 5.5 Accuracy recovery with continued pretraining
This section demonstrates the capability of the MoNE framework with the 25% pruned OLMoE after continued pretraining. The results are summarized in Table 2. This table shows that MoNE achieves the average accuracy closest to the original model with only 2B tokens from a pretraining dataset, demonstrating the promising capability of the MoNE computation framework. Besides, MC-SMoE and RS reclaim 8.54 and 14.02 average accuracy, indicating that expert pruning is not only effective to eliminate redundancy, but also relatively easier to recover performance with pretraining.
5.6 Inference latency and memory footprint
Figure 5: Inference latency and memory footprint over downstream tasks for the three models with different pruning ratios and methods. Latency are normalized with original models. Bars are normalized latency and lines are maximum allocated memory.
This section evaluates the inference latency and memory footprint for pruned models. For each model, we used the original HuggingFace inference implementation. Experiments were conducted on a single H20 GPU, using the first 20 samples from each downstream task with batch size 1 to eliminate padding effects. The maximum generation length was fixed to 256. Though different calibration data might affect the performance, no such variation was observed under this setup.
Figure 5 depicts the normalized latency and maximum allocated memory for models pruned by different methods and pruning ratios. Since the batch size is 1 and generation length is short, the main memory overhead is from model parameters. With the same pruning ratio, different methods achieve similar peak memory reduction. In contrast, different structured pruning methods can yield varying generation latency for MoE models. Angular prunes layers, thus achieving latency reduction linear to pruning ratio. RS and MC-SMoE show descent speedup on OLMoE model but cannot accelerate other two models. This may because the two models need to move routing scores to CPU to decide the expert execution, which downgrades the GPU utilization. Notably, MoE models pruned by FLAP cannot show any speedup. This may be attributed to the tile-quantization effect from the small weight matrices with irregular shapes within each expert [1]. MoNE shows speedup around 10% over original models, suggesting that a more efficient inference implementation is necessary to fully unlock the computational advantages of MoNE. We consider this as promising future work.
@@ 6 Conclusion
In this paper, we proposes MoNE, a novel expert pruning method that replaces redundant experts with lightweight novices to compress MoE models. MoNE evaluates the expert redundancy based on the access frequency and output variance of each expert within a transformer layer. Experts with low usage and stable output are pruned and replaced by novices, which are the unbiased estimation of their outputs to compress the model size while incurring minimal performance degradation. Extensive experiments on different MoE model architectures, calibration data sources and calibration sample sizes reveal that MoNE exceeds existing structured pruning methods by maintaining higher zero shot performance across nine downstream tasks, demonstrating the effectiveness and robustness of MoNE.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A More detailed results
@@
Table 3: Zero shot performance with 500 calibration samples from Zyda2 dataset.
@@
MoNE 36.69 66.04 73.36 86.00 41.07 35.60 75.73 58.12 69.69 60.26
Table 4: Zero shot performance with 1000 calibration samples from Zyda2 dataset.
@@
MoNE 37.71 65.36 73.49 84.00 41.22 36.60 75.30 57.40 69.53 60.07
Table 5: Zero shot performance with 100 calibration samples from C4 dataset.
@@
MoNE 35.24 55.47 68.20 87.00 22.94 35.40 76.50 53.43 68.35 55.84
Table 6: Zero shot performance with 500 calibration samples from C4 dataset.
@@
MoNE 34.39 55.72 67.68 84.00 22.90 35.60 76.28 54.51 67.72 55.42
Table 7: Zero shot performance with 1000 calibration samples from C4 dataset.
@@
MoNE 34.04 55.51 68.26 85.00 22.92 36.40 76.44 54.15 66.61 55.48
This section presents the experiment results on Zyda2 dataset with 500 and 1000 samples in Table 3 and Table 4. Table 5, Table 6 and Table 7 presents the experiment results on C4 dataset with 100, 500 and 1000 samples. The observations are the similar to those in Section 5.2 and 5.3.
Appendix B Redundant expert visualization
@@
This section complements the visualization of redundant experts for different models across the downstream tasks. The results are depicted in Figure 6 - Figure 13. As mentioned in Figure 1, for each figure, expert in blue circles has both high frequency and variance. Expert in red circles only has high variance. Expert in green circles only has high frequency. For each model across the nine downstream tasks, we can always identify the same important experts, validating the effectiveness of the redundancy metric, i.e., the expert access frequency and output variance.
Appendix C Comprehensive ablation study results
@@
Appendix D Limitations
The proposed MoNE is an expert pruning method for the structured pruning of MoE models and the experiments demonstrate excellent performance. However, there are a few limitations that are not covered in this paper. First, we do not conduct experiments on larger MoE models. The largest MoE models employed in this paper are at the 16B scale. Larger models may require more redundancy features other than the two metrics employed in this paper. However, those large models also need more computation resources for experiments. Second, we do not explore the combination of different structured pruning methods. While this paper focuses on expert pruning which is unique to MoE models and demonstrates the advantages of expert pruning over other structured pruning methods, combining different pruning methods can potentially expand the search space, leaving finer granularity to retain more model capabilities. We leave the two avenues as promising directions for future work.

### v2

@@ 5.3 Robustness evaluation
This section evaluates the robustness of MoNE across three key dimensions: model architecture, calibration data source, and calibration sample size. Due to prohibitive compute for exhaustive experiments on large models, we tested three models: OLMoE, Moonlight and Deepseek-V2-Lite using 100, 500 and 1000 calibration samples on C4 and Zyda-2 dataset separately. For each dimension, we vary one factor while averaging results over the other two, measuring both average accuracy and the standard deviation of accuracy drop. The results are visualized in Figure 3, with detailed scores provided in Appendix C. As shown in Figure 3(a), MoNE advances the Pareto frontier across all three dimensions at the 25% pruning ratio, demonstrating superior robustness and effectiveness compared to existing structured pruning methods. At the 50% pruning ratio (Figure 3(b)), MoNE exhibits slightly higher variance under varying model architectures and calibration sample sizes. Nevertheless, it remains the most effective method, outperforming baseline methods by a significant margin of 2.85.
@@ 5.4 Ablation study
Figure 4: Ablation study on expert access frequency, output variance and novice replacement. Numbers are the difference to the proposed MoNE. The detailed result is provided in Appendix E.
This section presents the ablation study to evaluate the effects of the two redundancy metrics and the impact of novice replacement across the downstream tasks. Figure 4 displays the average accuracy drop relative to our proposed methods, with lower values indicating greater degradation. Results are averaged over the three evaluation dimensions to provide a robust assessment. We observe that integrating the fused expert redundancy score with novice replacement yields better performance, particularly under higher pruning ratios. This indicates that our approach is especially effective in preserving model quality when pruning is more aggressive. Notably, for tasks such as BoolQ, COPA, and PIQA, our proposed method outperforms the ablation baselines by a large margin—achieving accuracy gains of up to 8.85. However, for MMLU, pruning based solely on frequency appears to offer a slight advantage, suggesting that frequently activated experts may play a more critical role in domain-specific reasoning tasks.
@@ 5.5 Accuracy recovery with continued pretraining
To evaluate performance recovery capabilities, we conducted continued pretraining on the 25% compressed OLMoE model pruned by 100 Zyda2 samples, as only this model releases its pretraining dataset, OLMoE-mix-0924222https://huggingface.co/datasets/allenai/OLMoE-mix-0924. The sequence length was set to 4096 and the global token size per step was 4M. Each pruned model was trained with 2B tokens, i.e., 512 steps, and the peak and minimum learning rate (lr) were 5e-5 and 5e-6, respectively. We employed the cosine lr scheduler with 50 warm up steps. Other hyperparameters were the same as the original configuration for OLMoE (Muennighoff et al., 2025). All the experiments could run on a single H20 GPU, but we accelerated the training with 16 H20 GPUs.
The results are summarized in Table 3. This table shows that MoNE achieves the average accuracy closest to the original model with only 2B tokens from a pretraining dataset, demonstrating the promising capability of the MoNE computation framework. Besides, MC-SMoE and RS reclaim 8.54 and 14.02 average accuracy, indicating that expert pruning is not only effective to eliminate redundancy, but also relatively easier to recover performance with continued pretraining.
@@ 6 Conclusion
In this paper, we propose MoNE, a novel expert pruning method that replaces redundant experts with lightweight novices to compress MoE models. MoNE evaluates expert redundancy based on expert access frequency and output variance in each model layer, pruning experts with low usage and stable outputs while replacing them with novices that provide unbiased output estimates. Extensive experiments across different MoE architectures, calibration data sources, and sample sizes demonstrate that MoNE outperforms existing structured pruning methods by maintaining higher zero-shot performance across nine downstream tasks.
Acknowledgment
We would like to acknowledge that computational work involved in this research work is supported by NUS IT’s Research Computing group using grant numbers CFP02-CF-004. Yang You’s research group is being sponsored by NUS startup grant (Presidential Young Professorship), Singapore MOE Tier-1 grant, ByteDance grant, NUS ARTIC grant, Apple grant, Alibaba grant, Google Research and Google grant for TPU usage.
Appendix A Specialized task evaluation
Table 4: Base and Instruct OLMoE models on Math and GSM8K.
Model GSM8K Math
No calibration
OLMoE 52.77 15.78
OLMoE-Instruct 67.63 18.64
Calibration: Zyda2
OLMoE 3.34 1.94
OLMoE-Instruct 7.96 2.54
Calibration: Task-specific
OLMoE-Instruct (Math) 65.28 18.22
OLMoE-Instruct (GSM8K) 67.48 17.64
Table 5: Base and Instruct Moonlight models on Math and GSM8K.
Model GSM8K Math
No calibration
Moonlight 74.22 42.32
Moonlight-Instruct 77.03 39.26
Calibration: Zyda2
Moonlight 46.40 3.32
Moonlight-Instruct 51.86 6.64
Calibration: Task-specific
Moonlight-Instruct (Math) 75.28 38.56
Moonlight-Instruct (GSM8K) 76.72 35.84
Table 6: Base and Instruct DeepSeek-V2-Lite models on Math and GSM8K.
Model GSM8K Math
No calibration
DeepSeek-V2-Lite 38.82 16.54
DeepSeek-V2-Lite-Chat 66.49 18.56
Calibration: Zyda2
DeepSeek-V2-Lite 25.40 6.80
DeepSeek-V2-Lite-Chat 37.07 4.76
Calibration: Task-specific
DeepSeek-V2-Lite-Chat (Math) 64.44 20.76
DeepSeek-V2-Lite-Chat (GSM8K) 63.76 19.70
Table 7: Base and Instruct Qwen2-57B-A14B models on Math and GSM8K.
Model GSM8K Math
No calibration
Qwen2-57B-A14B 79.68 40.50
Qwen2-57B-A14B-Instruct 69.90 31.30
Calibration: Zyda2
Qwen2-57B-A14B 74.07 31.28
Qwen2-57B-A14B-Instruct 64.90 22.84
Calibration: Task-specific
Qwen2-57B-A14B-Instruct (Math) 58.45 29.56
Qwen2-57B-A14B-Instruct (GSM8K) 52.92 29.62
Table 8: Base and Instruct Qwen3-30B-A3B models on Math and GSM8K.
Model GSM8K Math
No calibration
Qwen3-30B-A3B 85.37 50.48
Qwen3-30B-A3B-Instruct 90.90 46.94
Calibration: Zyda2
Qwen3-30B-A3B 71.72 9.44
Qwen3-30B-A3B-Instruct 78.01 9.88
Calibration: Task-specific
Qwen3-30B-A3B-Instruct (Math) 89.76 48.52
Qwen3-30B-A3B-Instruct (GSM8K) 90.30 51.70
We extended MoNE to two specialized tasks: Math and GSM8K. We conducted experiments on both base models and instruct models. To calibrate the instruct models, we adopted the first 100 samples from the training dataset of each task. The results are reported in Tables 48.
Our experiments on specialized tasks reveal two important findings regarding the effectiveness of different calibration strategies. After pruning with pretraining data (Zyda2), larger base models preserve significantly more accuracy, which is consistent with our observations from general tasks in Section 5.2. However, pretraining data cannot accurately capture the distribution of specialized tasks, leading to accuracy drops of up to 49% for the smallest model, OLMoE. This demonstrates that while model scale provides some resilience to pruning, the mismatch between pretraining data and task-specific distributions poses substantial challenges for maintaining performance on specialized tasks.
In contrast, MoNE can effectively preserve performance by pruning instruct models with training data from the specialized tasks themselves. For example, OLMoE-Instruct incurs only minimal accuracy drops of at most 1% for both tasks when calibrated with task-specific data. This result is not surprising, as state-of-the-art models also rely on specialized fine-tuning approaches such as supervised fine-tuning and reinforcement learning with domain-specific data to enhance their capability in these tasks (Muennighoff et al., 2025; Liu et al., 2025; Qwen, 2024; Yang et al., 2025). The effectiveness of task-specific calibration suggests that the distribution alignment between calibration data and target tasks plays a critical role in determining post-pruning performance.
In summary, we acknowledge that calibration via pretraining data is insufficient to preserve model capability for specialized tasks, and task-specific calibration can effectively mitigate this issue. A promising future direction for model compression is to develop methods that bridge the gap between these two calibration approaches, potentially through hybrid calibration strategies or adaptive data selection techniques that better capture task-specific distributions while maintaining broad coverage.
Appendix B Pruning metric ablation study
We conducted an extended ablation study on the redundancy score to evaluate the sensitivity of MoNE to different scoring formulations. Following the same evaluation configuration as in Section 5.2, we compared MoNE against five variants: normalized scores, log-sum aggregation, and weighted sum with varying emphasis on output variance (25%, 50%, and 75%). The results are reported in Table 9.
Table 9: Ablation study on redundancy scoring methods across different models. All values represent average performance scores.
Model MoNE Normalized Log-sum Weighted 25% Weighted 50% Weighted 75%
OLMoE 61.04 61.39 61.23 57.07 57.07 57.07
Moonlight 69.10 68.58 69.05 68.95 69.03 69.07
DeepSeek-V2-Lite 66.12 65.70 66.02 65.63 65.66 66.07
Qwen2-57B-A14B 71.75 71.73 71.75 72.55 72.57 72.53
Qwen3-30B-A3B 72.69 73.62 73.66 71.96 72.29 72.48
The results indicate that the pruning score exhibits minimal scale sensitivity across different formulations. In particular, the log-sum aggregation shows nearly identical results to MoNE, which is expected since $\log(\theta^{var})+\log(\theta^{freq})=\log(\theta^{var}\times\theta^{freq})$ does not affect the partial ordering during expert redundancy ranking. This mathematical equivalence ensures that both formulations produce the same pruning decisions despite their different representations.
We did not include learnable weights in this ablation study, as training five models with learnable parameters would be prohibitively expensive and time-consuming. However, the weighted sum results reveal model-dependent behavior, where only Qwen2-57B-A14B achieves minor improvements of approximately 0.8% with weighted scoring. In contrast, smaller models like OLMoE incur severe performance drops of around 4% under the weighted sum formulation. This suggests that while larger models may benefit from adjusting the relative importance of variance and frequency components, smaller models are more sensitive to such modifications and perform better with the balanced geometric mean approach used in MoNE.
Appendix C More detailed results
@@
Table 10: Zero shot performance with 100 calibration samples from Zyda2 dataset. Best results are in bold, and the second best are underlined.
0% OLMoE 49.23 76.89 70.09 85.00 53.54 44.40 79.76 71.84 68.90 66.63
FLAP 29.18 54.92 62.17 68.00 30.51 29.60 67.57 55.23 56.27 50.39
MC-SMoE 24.49 31.44 59.33 67.00 23.01 26.00 53.92 53.43 53.12 43.53
RS 21.50 28.62 39.45 61.00 23.27 26.00 52.34 51.99 51.85 39.56
MoNE (Ours) 28.16 40.24 63.12 78.00 25.21 32.40 63.33 60.65 63.54 50.52
FLAP 33.87 61.36 63.30 75.00 36.80 36.00 69.37 57.04 62.12 54.98
MC-SMoE 29.52 47.94 59.54 79.00 23.94 31.40 67.30 57.04 60.46 50.68
RS 37.80 58.42 70.86 89.00 23.27 38.00 78.18 57.76 70.80 58.23
MoNE (Ours) 43.09 70.03 76.12 90.00 23.57 40.80 78.78 58.84 70.17 61.27
FLAP 35.24 60.31 69.66 79.00 36.13 35.20 74.76 56.68 64.09 56.79
MC-SMoE 24.57 35.82 56.36 67.00 28.29 27.40 61.37 52.35 53.91 45.23
RS 36.01 57.45 57.98 89.00 24.91 40.80 78.02 54.15 62.75 55.67
MoNE (Ours) 37.20 67.17 73.39 84.00 42.30 36.80 75.30 59.57 67.88 60.40
0% Qwen2-57B-A14B 49.66 69.44 86.45 93.00 74.06 44.20 81.23 74.73 74.27 71.89
50% Angular 25.17 31.19 45.08 54.00 23.74 29.80 54.62 53.07 50.28 40.77
FLAP 36.35 61.95 73.43 80.00 46.50 38.20 74.59 68.95 68.59 60.95
MC-SMoE 41.64 64.73 79.91 87.00 59.65 38.80 77.53 71.84 68.27 65.49
RS 20.90 29.38 50.83 57.00 23.39 25.40 52.88 51.26 49.25 40.03
MoNE (Ours) 45.43 65.22 80.99 85.00 62.00 42.00 78.41 70.90 69.51 66.61
0% Qwen3-30B-A3B 55.89 79.42 88.69 84.00 77.82 44.80 80.30 82.31 70.64 73.76
50% Angular 26.37 39.48 50.86 59.00 24.65 28.80 60.17 55.23 48.46 43.67
FLAP 36.52 61.99 70.80 78.00 48.92 34.00 70.57 74.37 64.96 60.01
MC-SMoE 39.76 63.09 85.84 83.00 56.30 39.40 74.54 78.34 69.14 65.49
RS 48.04 73.06 86.61 83.00 61.37 42.40 79.43 70.40 68.27 68.06
MoNE (Ours) 44.62 67.05 85.41 85.00 54.58 41.60 79.00 71.84 67.80 66.32
Table 11: Zero shot performance with 500 calibration samples from Zyda2 dataset. Best results are in bold, and the second best are underlined.
@@
MoNE (Ours) 36.69 66.04 73.36 86.00 41.07 35.60 75.73 58.12 69.69 60.26
Table 12: Zero shot performance with 1000 calibration samples from Zyda2 dataset. Best results are in bold, and the second best are underlined.
@@
MoNE (Ours) 37.71 65.36 73.49 84.00 41.22 36.60 75.30 57.40 69.53 60.07
Table 13: Zero shot performance with 100 calibration samples from C4 dataset. Best results are in bold, and the second best are underlined.
@@
MoNE (Ours) 35.24 55.47 68.20 87.00 22.94 35.40 76.50 53.43 68.35 55.84
Table 14: Zero shot performance with 500 calibration samples from C4 dataset. Best results are in bold, and the second best are underlined.
@@
MoNE (Ours) 34.39 55.72 67.68 84.00 22.90 35.60 76.28 54.51 67.72 55.42
Table 15: Zero shot performance with 1000 calibration samples from C4 dataset. Best results are in bold, and the second best are underlined.
@@
MoNE (Ours) 34.04 55.51 68.26 85.00 22.92 36.40 76.44 54.15 66.61 55.48
This section presents the experiment results on Zyda2 dataset with 100, 500 and 1000 calibration samples in Table 10, Table 11 and Table 12. Table 13, Table 14 and Table 15 presents the experiment results on C4 dataset with 100, 500 and 1000 calibration samples. The observations are similar to those in Section 5.2 and Section 5.3. Nevertheless, all methods incurs more aggressive performance drop at higher (50%) pruning ratio. Such phenomenon is inevitable for all model compression approaches and inspires the development of MoNE to push the performance limits of structured pruning.
Appendix D Redundant expert visualization
@@
This section complements the visualization of redundant experts for OLMoE, Moonlight and Deepseek-V2-Lite across the nine downstream tasks. The results are depicted in Figure 5 - Figure 12. As mentioned in Figure 1, for each figure, expert in blue circles has both high frequency and variance. Expert in red circles only has high variance. Expert in green circles only has high frequency. For each model across the nine downstream tasks, we can always identify the same important experts, validating the effectiveness of the redundancy metric, i.e., the expert access frequency and output variance.
Appendix E Comprehensive ablation study results
@@
Appendix F Inference latency and memory footprint
Table 16: Inference latency and memory footprint under different pruning ratio, input sequence and batch size.
Pruning ratio Seed Batch size Novice hit ratios* Total latency (s) Speedup Memory usage (GB)
0% 1020 1 0 28.86 1 62.72
25% 1020 1 0.08 27.84 1.04 47.72
50% 1020 1 0.35 26.27 1.10 32.73
0% 1998 1 0 29.01 1 62.72
25% 1998 1 0.06 28.61 1.01 47.72
50% 1998 1 0.34 26.22 1.11 32.73
0% 1020 128 0 183.12 1 66.00
25% 1020 128 0.08 174.60 1.05 51.00
50% 1020 128 0.34 141.07 1.30 36.01
0% 1998 128 0 182.37 1 66.00
25% 1998 128 0.08 169.66 1.07 51.00
50% 1998 128 0.33 142.38 1.28 36.01
0% 1020 512 0 248.02 1 75.90
25% 1020 512 0.09 212.40 1.17 60.90
50% 1020 512 0.33 189.57 1.31 45.91
0% 1998 512 0 245.42 1 75.90
25% 1998 512 0.08 212.88 1.15 60.90
50% 1998 512 0.34 180.92 1.36 45.91
*Novice hit ratios: the portion of tokens routed to novices across the model.
While this work mainly focuses on enhancing performance preserving capability of structured pruning, this section evaluates the inference latency and memory footprint for pruned models. We integrated the pruned Qwen3-30B-A3B with the popular inference framework, SGLang with transformers backend. We fixed the 512 input tokens and 256 output tokens, and profiled the latency and memory footprint with SGLang built-in utilities. We varied the random seed to generate different input sequences. Experiments were conducted on a single H20 GPU. The results are listed in Table 16.
According to Table 16, we would like to clarify three points regarding the latency and memory usage. First, inference latency speedup of MoNE is sensitive to batch size. When batch size is 1, there is minor speedup, as the decoding phase only executes one tokens per step, leading to a memory bound situation where the GPU tensor core is underutilized and the major bottleneck is the GPU memory bandwidth (Hong et al., 2023; Frantar et al., 2025). By increasing the batch size to 512, MoNE can achieve speedup of 36%. Second, inference latency speedup is sensitive to novice hit ratio instead of novice counts (pruning ratio). Qwen3-30B-A3B has 128 experts per layer, and each input token is routed to 8 experts per novices. The actual speedup is decided by the routing results for each input token at runtime. With higher novice hit ratio, MoNE is expected to attain more speedup. While higher pruning ratio may lead to more novices in each layer, it cannot directly translate to speedup. For the most extreme case where there is no token routed to novices, the computation overhead (as well as the accuracy) will be the same as the original model. Finally, Table 16 indicates that MoNE achieves maximum memory reduction that consistently increases with the pruning ratio. For a given pruning ratio, the heavy expert parameters are directly replaced by a light weight constant vector.